A rocket rises vertically from rest with an acceleration of 3.2 m/s2 until it runs out of fuel at 775 m. After this point, its acceleration is only that of gravity, downward. (a) What is the velocity of the rocket when it runs out of fuel? (b) How long does it take to reach this point? (c) What maximum altitude does the rocket reach? (d) How much time does it take (total) to reach its maximum altitude? (e) With what velocity does it strike the earth? (f) How long (total) is it in the air?
Answer #1

(a) v^2-u^2=2as

v^2-0= 2as

v=(2as)1/2 =(2x3.2x775)1/2

v = 70.4 m/s

(b) v-u=at

t=v/a = 70.4/3.2

t=22.08 s

(c) v^2-u^2= 2as

0-v^2= -2gh

h=v^2/2g = 70.4x70.4/2x9.8

h = 252.86 m

H =h+s = 252.86+775

H=1027.86 m

(d) v-u=at

0-v= -gt

t=v/g =70.4/9.8=7.18 s

ta = t in(b) +t= 22.08+7.18

Time of asending ta =29.26 s

(e) v^2-u^2=2gH

v =(2gH)^1/2 =(2x9.8x1027.86)^1/2

v = 141.93 m/s

(f) time of desendin td = v/g = 141.93/9.8 =14.48 s

Total time of filght T =ta+td = 29.26+14.48

T = 43.74 s